OpenStudy (anonymous):

s = 2 sin πt + 2 cos πt

OpenStudy (dumbcow):

are you looking for the graph? http://www.wolframalpha.com/input/?i=plot+2sin%28pi*t%29+%2B2cos%28pi*t%29+for+t+%3D+0+to+2

OpenStudy (dumbcow):

it can also be written as \[2\sqrt{2} \sin(\frac{\pi}{4} + \pi t)\]